a certain type of gutter comes in 6 foot and 10 foot sections. How many different lengths can be formed usin g3 sections of gutter?

Found 2 solutions by Fombitz, JBarnum:
Answer by Fombitz(32388)

6, 6, 6=18
6, 6, 10=22
6, 10, 6=22
6, 10, 10=26
10, 6, 6=22
10, 6, 10=26
10, 10, 6=26
10, 10, 10=30
8 (2^3) different configurations yield 4 different lengths (18,22,26,30).

Answer by JBarnum(2146)

well if u write out the 4 possibilities
6+6+6=18
6+6+10=22
6+10+10=26
10+10+10=30
so there are only 4 different lengths